# CAST: Component-Aligned 3D Scene Reconstruction from an RGB Image

### v1

@@ 4.3 Iterative Generation Procedure
$\bm{z}^{(k)}=\text{ObjectGen}(\bm{c},\bm{p}^{(k)}\otimes\beta^{(k)}).$ (6)
@@
Once the object geometry is determined, we apply a state-of-the-art texture generation module to create photo-realistic surface details. Following established texture synthesis pipelines [90, 92], we assign UV mappings and train a generative network to paint detailed textures onto the 3D meshes. This module is designed to robustly handle images under various augmentation, ensuring that the final textures match the input appearance even under occlusion or limited visibility.
Figure 4: Physics-aware correction via constraint graph mapped from fine-grained relation graph. Top: Floating surfboard grounded on the van. Bottom: Penetrating guitar and cooler separated.
Figure 5: Bringing the vibrant diversity of the real world into the virtual realm, this collection reimagines open-vocabulary scenes as immersive digital environments, capturing the richness and depth of each unique setting. For each scene, the images display as follows: the top-left shows the input image, the top-center displays the rendered geometry, and the right presents the rendered image with realistic textures.
5 Physics-Aware Correction
@@
To address these issues, we introduce a physics-aware correction process that optimizes the rotation and translation of objects, ensuring the scene adheres to physical constraints consistent with common sense. The correction process is motivated by physical simulation (Sec. 5.1) and formulated as an optimization problem (Sec. 5.2) based on inter-object relationships represented by a scene graph (Sec. 5.3) extracted from the image.
5.1 A Quick Primer to Rigid-Body Simulation
We introduce the fundamental principles of physical (rigid-body) simulation, which inspire our problem formulation and make our framework more accessible to downstream applications such as gaming and robotics. For a thorough survey, we refer the readers to [5].
In rigid-body simulations, the world is modeled as an ordinary differential equation (ODE) process. In each simulation step, it begins with the Newton-Euler (differential) equations, which describe the dynamic motion of rigid bodies in the absence of contact. Collision detection is conducted to find the contact points between rigid bodies, which are needed to determine contact forces. For contact handling and collision resolution, there are usually several conditions: non-penetration constraints to prevent bodies from overlapping, a friction model ensuring contact forces remain within their friction cones, and complementarity constraints that enforce specific disjunctive relationships among variables. Solvers are used to resolve the system comprising equations and inequalities, subsequently updating the velocity and position of each rigid body.
@@ 5.1 A Quick Primer to Rigid-Body Simulation
2) Imperfect Geometries: While our 3D generative model produces high-quality geometries, minor imperfections may still occur. Rigid-body simulators typically require convex decomposition [52, 53, 76] of objects, which introduces additional complexity and hyperparameters. Overly fine-grained decomposition can result in non-flat, complex surfaces, causing objects to fall or move unexpectedly during simulation. Conversely, coarse decomposition may lead to visually floating objects due to discrepancies between the visual and collision geometries.
@@
5.2 Problem Formulation and Physical Constraints
@@
$\min_{\mathcal{T}=\{T_{1},T_{2},\dots,T_{N}\}}\sum_{i,j}C(T_{i},T_{j};\bm{o}_{i},\bm{o}_{j})$ (8)
where $N$ is the number of objects, $T_{i}$ is the rigid transformation (rotation and translation) of the i-th object $\bm{o}_{i}$. $C$ is the cost function representing the relationship between $\bm{o}_{i}$ and $\bm{o}_{j}$. Note that the cost function varies depending on the type of relationship.
Motivated by physical simulation, we categorize the relationships into two types: contact and support. The relationships are identified with the assistance of a VLM, as detailed in Sec. 5.3.
@@
$\begin{split}C(T_{i},T_{j};\bm{o}_{i}\to\bm{o}_{j})=\frac{\sum_{p\in\partial\bm{o}_{j}}D_{i}(p(T_{j}))\mathbb{I}(D_{i}(p(T_{j}))<0)}{\sum_{p\in\partial\bm{o}_{j}}\mathbb{I}(D_{i}(p(T_{j}))<0)}\\ +\max(\min_{p\in\partial\bm{o}_{j}}D_{i}(p(T_{j})),0)\\ C(T_{i},T_{j};\bm{o}_{j}\to\bm{o}_{i})=\frac{\sum_{p\in\partial\bm{o}_{i}}D_{j}(p(T_{i}))\mathbb{I}(D_{j}(p(T_{i}))<0)}{\sum_{p\in\partial\bm{o}_{i}}\mathbb{I}(D_{j}(p(T_{i}))<0)}\\ +\max(\min_{p\in\partial\bm{o}_{i}}D_{j}(p(T_{i})),0)\\ C(T_{i},T_{j})=C(T_{i},T_{j};\bm{o}_{i}\to\bm{o}_{j})+C(T_{i},T_{j};\bm{o}_{j}\to\bm{o}_{i})\\ \text{if }\bm{o}_{i}\text{ and }\bm{o}_{j}\text{ are in contact}\end{split}$ (9)
@@
$C(T_{i},T_{j})=|\min_{p\in\partial\bm{o}_{j}}D_{i}(p(T_{j})|,\ \text{if }\bm{o}_{i}\text{ supports }\bm{o}_{j}$ (10)
@@
$C(T_{i},T_{j})=\frac{\sum_{p\in\partial\bm{o}_{j}}D_{i}(p(T_{j})\mathbb{I}(0<D_{i}(p)<\sigma)}{\sum_{p\in\partial\bm{o}_{j}}\mathbb{I}(0<D_{i}(p)<\sigma)}$ (11)
@@
5.3 Scene Relation Graph
Physical cues, particularly inter-object relationships, are visually present in the image. We leverage the strong common-sense reasoning capabilities [62, 42, 14] of visual-language models, specifically GPT-4v [1], to identify pairwise physical constraints as defined in Sec. 5.2. Given an image, we employ the Set of Mark [83] (SoM) technique to visually prompt GPT-4v to describe the inter-object relationships, and subsequently extract a scene relation graph from the answers. To address the sampling uncertainty inherent in VLMs, we adopt an ensemble strategy, combining results from multiple trials to produce a robust inferred graph.
Instead of directly asking GPT-4v to identify Support and Contact relationships, we first provide it with more fine-grained physical relationships, such as Stack (Object 2 supports Object 1), Lean (Object 1 leans against Object 2), and Hang (Object 2 supports Object 1 from above). We then map these detailed relationships to the predefined categories of Support and Contact for further optimization. Specifically, if there are edges pointing toward each other between two nodes, the edge is categorized as Contact; otherwise, it is categorized as Support. Prompting GPT-4v with these nuanced relationships helps eliminate potential ambiguity in binary relationship classification and facilitates more accurate reasoning by GPT-4v. An example of the resulting graph is illustrated in Fig. 4.
The mapped scene constraint graph is a directed graph where nodes represent object instances and edges denote physical relationships between objects. A Contact relationship is represented by a bidirectional edge, while a Support relationship is depicted as a directed edge. This graph serves as the foundation for defining the cost functions used in Eq. 8.
5.4 Optimization with Physics-Aware Relation Graph
@@
Figure 6: Qualitative comparisons of CAST with state-of-the-art single-image scene reconstruction methods. From left to right: Input image, CAST, ACDC, and Gen3DSR.
6 Result
@@ 6.1 Implementation Details
The ObjectGen (Sec. 4.1) model’s pretraining follows the methodology outlined in 3DShape2VecSet [90] and CLAY [92], where we leverage both a Variational Autoencoder (VAE) and a Latent Diffusion Model (LDM) to generate 3D object geometries. Both the VAE and LDM modules are implemented using a 24-layer transformer, comprising a total of 1.5 billion parameters. The model is trained on the Objaverse [20] dataset, which consists of approximately 500,000 3D assets after filtering. The partial point cloud conditioning follows a similar approach to CLAY’s adaptation framework. We encode the canonical-space partial point cloud as positional embeddings with a feature dimension of 512, which are injected into the main LDM transformer using cross-attention mechanisms. For each 3D asset, we render 32 views and precompute depth maps using MoGe [74] and Metric3D [86]. These depth maps are then lifted into point clouds during training, with random masks applied to simulate occlusions. We sample 2048 points from the unprojected point cloud using Farthest Point Sampling (FPS), which serves as conditioning input for the LDM. To enhance the model’s robustness, we randomly interpolate between the ground truth and predicted partial point clouds, allowing the system to handle data of varying quality. The conditioning module is trained on 200K curated data from Objaverse over 3000 epochs with 64 Nvidia A800 GPUs. The AdamW optimizer is used with a learning rate of 1e-5.
@@
The AlignGen (Sec. 4.2) module, responsible for generating pose alignment, utilizes a 24-layer transformer with a feature dimension of 512, resulting in a total of 150 million parameters. During training, we randomly sample a partial point cloud in the canonical space from the point clouds lifted from precomputed depth maps and apply a random transformation to this point cloud. The transformed point cloud, along with the geometry latent code $\bm{z}$ from ObjectGen, is used as the conditioning input. FPS is utilized to sample 2048 points from the partial point cloud to ensure a fixed number of inputs to the transformer. Training is conducted on the same 200K curated dataset over 1,500 epochs with 64 Nvidia A800 GPUs. The AdamW optimizer is used with a learning rate of 1e-5.
Table 1: Quantitative comparison of scene reconstruction methods across four metrics with CLIP score, GPT-4 ranking, user study of visual quality (VQ), and physical plausibility (PP).
@@
6.2 Comparison
Table 2: Quantitative comparison of scene reconstruction performance on the 3D-Front indoor dataset. We evaluate different methods based on Chamfer Distance (CD) for shape accuracy, F-Score (FS) for object-level reconstruction quality, and Intersection over Union (IoU) for scene-level overlap.
@@ 6.2 Comparison
We first evaluate our method, CAST, against state-of-the-art single-image scene reconstruction techniques on open-vocabulary scenarios. Fig. 6 illustrates the performance of three methods—(1) the retrieval-based approach ACDC [18], (2) the generation-based method Gen3DSR [21], and (3) our proposed CAST—across both reference and novel views. Our results highlight CAST’s superior ability to accurately reconstruct scenes in diverse settings, including indoor and outdoor environments, close-up perspectives, and AI-generated imagery. Existing retrieval-based methods (e.g., ACDC) often rely on fixed object categories (such as furniture) and thus offer limited flexibility. Meanwhile, generation-based approaches like Gen3DSR struggle with challenging geometries and partial occlusions, showing a rapid decrease in performance when rendering novel views. By contrast, CAST demonstrates robust scene reconstructions under varying conditions, underscoring its versatility for a broad range of real-world and generated scenarios.
To assess both the visual fidelity and semantic accuracy of the generated scenes, we employ two complementary evaluation methods including CLIP Score [93] and GPT-4 Reasoning. We compute the CLIP score between the rendered scene and the input image to measure overall reconstruction quality and visual similarity. To minimize environmental distractions, we remove backgrounds from both the rendered and reference images before computing the score. We additionally leverage GPT-4 to rank the generated scenes based on various semantic aspects, including object arrangement, physical relationships, and scene realism. This semantic feedback helps identify alignment or contextual errors that might not be apparent through pixel-based scores alone.
@@
Figure 7: We evaluate the generation performance with and without the occlusion-aware generation module. The RGB and normal renderings of the object highlight the significance of this module in ensuring the completeness and high quality of the generated object.
@@
Although CAST is designed to handle open-vocabulary scenes, many such scenes lack mesh ground truth, which complicates direct quantitative comparisons. To address this, we perform additional evaluations on the 3DFront dataset [22]. This dataset offers ground-truth meshes alongside corresponding rendered images, enabling a more precise assessment of both object-level and scene-level reconstructions. We compare our method with InstPIFu [45], ACDC [18], and Gen3DSR [21]. We compute Chamfer Distance and F-Score at the object level, as well as IoU, Chamfer Distance, and F-Score at the scene level, to assess both the fidelity of individual object geometries and the accuracy of their spatial layout. To ensure fairness, we replace the segmentation modules in other methods with ground-truth (GT) masks so that any differences stem purely from reconstruction ability rather than object partitioning.
@@
6.3 Evaluations
Figure 8: A stack of books with varying lengths and widths directly generated as a single complex object, demonstrating how point cloud conditioning enhances the preservation of scale, dimensions, and local details compared to traditional methods.
Figure 9: Comparative evaluation of pose estimation methods. Our pose alignment module demonstrates superior alignment accuracy compared to Iterative Closest Point (ICP) and differentiable rendering (DR).

### v2

@@ 4.3. Iterative Generation Procedure
(6) $\bm{z}^{(k)}=\text{ObjectGen}(\bm{c},\bm{p}^{(k)}\otimes\beta^{(k)}).$
@@
Once the object geometry is determined, we apply a state-of-the-art texture generation module to create photo-realistic surface details. Following established texture synthesis pipelines (Zhang et al., 2023, 2024a), we assign UV mappings and train a generative network to paint detailed textures onto the 3D meshes. This module is designed to robustly handle images under various augmentation, ensuring that the final textures match the input appearance even under occlusion or limited visibility.
Figure 4. Physics-aware correction via constraint graph mapped from fine-grained relation graph. Top: Floating surfboard grounded on the van. Bottom: Penetrating guitar and cooler separated.
Figure 5. Bringing the vibrant diversity of the real world into the virtual realm, this collection reimagines open-vocabulary scenes as immersive digital environments, capturing the richness and depth of each unique setting. For each scene, the images display as follows: the top-left shows the input image, the top-center displays the rendered geometry, and the right presents the rendered image with realistic textures.
5. Physics-Aware Correction
@@
To address these issues, we introduce a physics-aware correction process that optimizes the rotation and translation of objects, ensuring the scene adheres to physical constraints consistent with common sense. The correction process is motivated by physical simulation (Sec. 5.1) and formulated as an optimization problem (Sec. 5.2) based on inter-object relations represented by a scene graph (Sec. 5.3) extracted from the image.
5.1. A Quick Primer to Rigid-Body Simulation
We introduce the fundamental principles of physical (rigid-body) simulation, which inspire our problem formulation and make our framework more accessible to downstream applications such as gaming and robotics. For a thorough survey, we refer the readers to Bender et al. (2012).
In rigid-body simulations, the world is modeled as an ordinary differential equation (ODE) process. In each simulation step, it begins with the Newton-Euler (differential) equations, which describe the dynamic motion of rigid bodies in the absence of contact. Collision detection is conducted to find the contact points between rigid bodies, which are needed to determine contact forces. For contact handling and collision resolution, there are usually several conditions: non-penetration constraints to prevent bodies from overlapping, a friction model ensuring contact forces remain within their friction cones, and complementarity constraints that enforce specific disjunctive relations among variables. Solvers are used to resolve the system comprising equations and inequalities, subsequently updating the velocity and position of each rigid body.
@@ 5.1. A Quick Primer to Rigid-Body Simulation
2) Imperfect Geometries: While our 3D generative model produces high-quality geometries, minor imperfections may still occur. Rigid-body simulators typically require convex decomposition (Mamou and Ghorbel, 2009; Mamou et al., 2016; Wei et al., 2022) of objects, which introduces additional complexity and hyperparameters. Overly fine-grained decomposition can result in non-flat, complex surfaces, causing objects to fall or move unexpectedly during simulation. Conversely, coarse decomposition may lead to visually floating objects due to discrepancies between the visual and collision geometries.
@@
5.2. Problem Formulation and Physical Constraints
@@
(8) $\min_{\mathcal{T}=\{T_{1},T_{2},\dots,T_{N}\}}\sum_{i,j}C(T_{i},T_{j};\bm{o}_{i},\bm{o}_{j})$
where $N$ is the number of objects, $T_{i}$ is the rigid transformation (rotation and translation) of the i-th object $\bm{o}_{i}$. $C$ is the cost function representing the relation between $\bm{o}_{i}$ and $\bm{o}_{j}$. Note that the cost function varies depending on the type of relation.
Motivated by physical simulation, we categorize the relations into two types: contact and support. The relations are identified with the assistance of a VLM, as detailed in Sec. 5.3.
@@
(9) $\begin{split}C(T_{i},T_{j};\bm{o}_{i}\to\bm{o}_{j})=-\frac{\sum_{p\in\partial\bm{o}_{j}}D_{i}(p(T_{j}))\mathbb{I}(D_{i}(p(T_{j}))<0)}{\sum_{p\in\partial\bm{o}_{j}}\mathbb{I}(D_{i}(p(T_{j}))<0)}\\ +\max(\min_{p\in\partial\bm{o}_{j}}D_{i}(p(T_{j})),0)\\ C(T_{i},T_{j};\bm{o}_{j}\to\bm{o}_{i})=-\frac{\sum_{p\in\partial\bm{o}_{i}}D_{j}(p(T_{i}))\mathbb{I}(D_{j}(p(T_{i}))<0)}{\sum_{p\in\partial\bm{o}_{i}}\mathbb{I}(D_{j}(p(T_{i}))<0)}\\ +\max(\min_{p\in\partial\bm{o}_{i}}D_{j}(p(T_{i})),0)\\ C(T_{i},T_{j})=C(T_{i},T_{j};\bm{o}_{i}\to\bm{o}_{j})+C(T_{i},T_{j};\bm{o}_{j}\to\bm{o}_{i})\\ \text{if }\bm{o}_{i}\text{ and }\bm{o}_{j}\text{ are in contact}\end{split}$
@@
(10) $C(T_{i},T_{j})=|\min_{p\in\partial\bm{o}_{j}}D_{i}(p(T_{j}))|,\ \text{if }\bm{o}_{i}\text{ supports }\bm{o}_{j}$
@@
(11) $C(T_{i},T_{j})=\frac{\sum_{p\in\partial\bm{o}_{j}}D_{i}(p(T_{j})\mathbb{I}(0<D_{i}(p)<\sigma)}{\sum_{p\in\partial\bm{o}_{j}}\mathbb{I}(0<D_{i}(p)<\sigma)}$
@@
5.3. Scene Relation Graph
Physical cues, particularly inter-object relations, are visually present in the image. We leverage the strong common-sense reasoning capabilities (Rana et al., 2023; Li et al., 2024b; Cheng et al., 2024) of visual-language models, specifically GPT-4v (Achiam et al., 2023), to identify pairwise physical constraints as defined in Sec. 5.2. Given an image, we employ the Set of Mark (Yang et al., 2023) (SoM) technique to visually prompt GPT-4v to describe the inter-object relations, and subsequently extract a scene relation graph from the answers. To address the sampling uncertainty inherent in VLMs, we adopt an ensemble strategy, combining results from multiple trials. We define relations as correct if they appear in more than half of the samples to produce a robust inferred graph. To be more specific, we apply the Set-of-Mark method with random colorization and numerical ordering multiple times,enabling more reliable and consistent outputs for further GPT-based question-answering tasks.
Instead of directly asking GPT-4v to identify Support and Contact relations, we first provide it with more fine-grained physical relations, such as Stack (Object 2 supports Object 1), Lean (Object 1 leans against Object 2), and Hang (Object 2 supports Object 1 from above). We instruct GPT-4v to analyze numbered objects from the Set-of-Mark method and output all contact-based relations, covering six types: Stack, Lean, Hang, Clamped, Contained, and Edge/Point. The prompt specifies that only contacting objects have relations and defaults to Stack for ambiguous cases.
We then map these detailed relations to the predefined categories of Support and Contact for further optimization. Specifically, if there are edges pointing toward each other between two nodes, the edge is categorized as Contact; otherwise, it is categorized as Support. Prompting GPT-4v with these nuanced relations helps eliminate potential ambiguity in binary relation classification and facilitates more accurate reasoning by GPT-4v. An example of the resulting graph is illustrated in Fig. 4.
The mapped scene constraint graph is a directed graph where nodes represent object instances and edges denote physical relations between objects. A Contact relation is represented by a bidirectional edge, while a Support relation is depicted as a directed edge. This graph serves as the foundation for defining the cost functions used in Eq. 8.
5.4. Optimization with Physics-Aware Relation Graph
@@
Figure 6. Qualitative comparisons of CAST with state-of-the-art single-image scene reconstruction methods. From left to right: Input image, CAST, ACDC, and Gen3DSR. Top to bottom: random open vocabulary dataset (rows 1–3), Gen3DSR input (rows 4–5), ACDC input (rows 6–7).
6. Result
@@ 6.1. Implementation Details
The ObjectGen (Sec. 4.1) model’s pretraining follows the methodology outlined in 3DShape2VecSet (Zhang et al., 2023) and CLAY (Zhang et al., 2024a), where we leverage both a Variational Autoencoder (VAE) and a Latent Diffusion Model (LDM) to generate 3D object geometries. Both the VAE and LDM modules are implemented using a 24-layer transformer, comprising a total of 1.5 billion parameters. The model is trained on the Objaverse (Deitke et al., 2023) dataset, which consists of approximately 500,000 3D assets after filtering. The partial point cloud conditioning follows a similar approach to CLAY’s adaptation framework. We encode the canonical-space partial point cloud as positional embeddings with a feature dimension of 512, which are injected into the main LDM transformer using cross-attention mechanisms. For each 3D asset, we render 32 views and precompute depth maps using MoGe (Wang et al., 2024b) and Metric3D (Yin et al., 2023). These depth maps are then lifted into point clouds during training, with random masks applied to simulate occlusions. We sample 2048 points from the unprojected point cloud using Farthest Point Sampling (FPS), which serves as conditioning input for the LDM. To enhance the model’s robustness, we randomly interpolate between the ground truth and predicted partial point clouds, allowing the system to handle data of varying quality. The conditioning module is trained on 200K curated data from Objaverse over 3000 epochs with 64 Nvidia A800 GPUs required approximately one week. The AdamW optimizer is used with a learning rate of 1e-5. For inferencing a single object, object generation takes approximately 7 seconds, and texture generation takes approximately 10 seconds per object on an NVIDIA A6000 GPU.
@@
The AlignGen (Sec. 4.2) module, responsible for generating pose alignment, utilizes a 24-layer transformer with a feature dimension of 512, resulting in a total of 150 million parameters. During training, we randomly sample a partial point cloud in the canonical space from the point clouds lifted from precomputed depth maps and apply a random transformation to this point cloud. The transformed point cloud, along with the geometry latent code $\bm{z}$ from ObjectGen, is used as the conditioning input. FPS is utilized to sample 2048 points from the partial point cloud to ensure a fixed number of inputs to the transformer. Training is conducted on the same 200K curated dataset over 1,500 epochs with 64 Nvidia A800 GPUs for approximately two day. The AdamW optimizer is used with a learning rate of 1e-5. During inference, AlignGen module takes around 1 second for one object for pose generation.
Table 1. Quantitative comparison of scene reconstruction methods across four metrics with CLIP score, GPT-4 ranking, user study of visual quality (VQ), and physical plausibility (PP).
@@
6.2. Comparison
Table 2. Quantitative comparison of scene reconstruction performance on the 3D-Front indoor dataset. We evaluate different methods based on Chamfer Distance (CD) for shape accuracy, F-Score (FS) for object-level reconstruction quality, and Intersection over Union (IoU) for scene-level overlap.
@@ 6.2. Comparison
We first evaluate our method, CAST, against state-of-the-art single-image scene reconstruction techniques on open-vocabulary scenarios. We also included images used by ACDC and Gen3DSR to further demonstrate the scene reconstruction results of different methods. Fig. 6 illustrates the performance of three methods—(1) the retrieval-based approach ACDC (Dai et al., 2024), (2) the generation-based method Gen3DSR (Dogaru et al., 2024), and (3) our proposed CAST—across both reference and novel views. Our results highlight CAST’s superior ability to accurately reconstruct scenes in diverse settings, including indoor and outdoor environments, close-up perspectives, and AI-generated imagery.
As shown in Fig.6, CAST distinguishes itself from both ACDC and Gen3DSR through innovative advancements. Unlike ACDC, which is limited to indoor scenes and relies on large datasets for object retrieval, often producing objects similar to those in the scene rather than the objects themselves, CAST supports open-vocabulary generalization. This allows CAST to accurately reconstruct objects in varied and complex environments. While ACDC uses simple bounding boxes as proxies, CAST combines image-based physical priors with mesh optimization to effectively manage complex scenes. In contrast to Gen3DSR, CAST employs direct 3D generation via a Masked Autoencoder, eliminating the error-prone 2D inpainting step. This results in smoother meshes, significantly outperforming Gen3DSR in single-object generation quality, particularly in challenging scenes. Moreover, Gen3DSR’s lack of simulation often leads to issues such as interpenetration or floating objects, causing scenes to appear consistent only from the input viewpoint and degrading novel view rendering. CAST, by contrast, ensures robust scene consistency across perspectives. CAST demonstrates robust scene reconstructions under varying conditions, underscoring its versatility for a broad range of real-world and generated scenarios.
To assess both the visual fidelity and semantic accuracy of the generated scenes, we employ two complementary evaluation methods including CLIP Score (Zhengwentai, 2023) and GPT-4 Reasoning. We compute the CLIP score between the rendered scene and the input image to measure overall reconstruction quality and visual similarity. To minimize environmental distractions, we remove backgrounds from both the rendered and reference images before computing the score. We additionally leverage GPT-4 to rank the generated scenes based on various semantic aspects, including object arrangement, physical relations, and scene realism. This semantic feedback helps identify alignment or contextual errors that might not be apparent through pixel-based scores alone.
@@
Figure 7. We evaluate the generation performance with and without the occlusion-aware generation module. The RGB and normal renderings of the object highlight the significance of this module in ensuring the completeness and high quality of the generated object.
@@
Although CAST is designed to handle open-vocabulary scenes, many such scenes lack mesh ground truth, which complicates direct quantitative comparisons. To address this, we perform additional evaluations on the 3DFront dataset (Fu et al., 2021). This dataset offers ground-truth meshes alongside corresponding rendered images, enabling a more precise assessment of both object-level and scene-level reconstructions. We compare our method with InstPIFu (Liu et al., 2022), ACDC (Dai et al., 2024), and Gen3DSR (Dogaru et al., 2024). We compute Chamfer Distance and F-Score at the object level, as well as IoU, Chamfer Distance, and F-Score at the scene level, to assess both the fidelity of individual object geometries and the accuracy of their spatial layout. To ensure fairness, we replace the segmentation modules in other methods with ground-truth (GT) masks so that any differences stem purely from reconstruction ability rather than object partitioning.
@@
6.3. Evaluations
Figure 8. A stack of books with varying lengths and widths directly generated as a single complex object, demonstrating how point cloud conditioning enhances the preservation of scale, dimensions, and local details compared to traditional methods.
Figure 9. Comparative evaluation of pose estimation methods. Our pose alignment module demonstrates superior alignment accuracy compared to Iterative Closest Point (ICP) and differentiable rendering (DR).